# 0Year 0Year 0Year

### 10 Mar 2026 v2

@@ 4.3 Main results
On BIOC-GS consisting of noisy clinical notes, we observe a clear trend across method categories: prompt-based methods outperform neural methods, which in turn surpass dictionary-based ones, consistent with prior findings (Groza et al., 2024a; Qi et al., 2024; Shlyk et al., 2024). This pattern reflects their respective modeling capabilities. Prompt-based methods use LLMs with strong language understanding abilities for entity linking, enabling them to better handle free-form and lexically diverse clinical text. Neural methods, typically fine-tuned on structured, ontology-aligned corpora, struggle to generalize to such noisy input. Dictionary-based methods perform worst, as they lack semantic understanding and rely solely on exact or approximate string matching.
@@
On GSC-2024 and ID-68, which feature more standardized text with less surface-level variation in entities, AutoPCR remains highly competitive, even though the advantage of language understanding becomes less pronounced. The top-performers on these two datasets, FastHPOCR and PhenoBERT, does not bring the advantage to the other; by contrast, AutoPCR consistently ranks behind them in the second place. This may be possibly because both methods partially annotated the datasets themselves, introducing inductive bias in their favor. Specifically, ID-68, as annotated by PhenoBERT, contains mentions that are not exactly matched by ontology concepts but instead linked to hypernyms (e.g., “thinning of the body of the corpus callosum” linked to HP:0002079 “Hypoplasia of the corpus callosum”). While this criterion may be helpful in certain use cases, it diverges from the standard task definition shared by our study and other competitive baselines. This inconsistency creates an ambiguous annotation standard: if every term covered by the ontology’s hierarchy is a potential candidate, the task essentially shifts into a classical hierarchical classification problem. Conversely, if only a subset is chosen, the labels become heavily influenced by the idiosyncratic styles of the annotators, making the problem more suitable for supervised learning rather than the zero-shot setting we address. This is also reflected in the experimental results, where PhenoBERT achieves an outlier precision on ID-68, hinting at a bias in the dataset’s construction. Nevertheless, AutoPCR still secures the second best, outperforming the next best baseline by approximately 3% in F1. Similarly, AutoPCR maintains a clear advantage over all other baselines except FastHPOCR on GSC-2024. This consistent lead demonstrates AutoPCR’s superior robustness, even when faced with deviating annotation criteria. Unlike on BIOC-GS, AutoPCR here significantly outperforms REAL, which can be credited to two novel designs: (1) a comprehensive entity extraction approach that improves recall on long, standardized text compared to LLM-based extraction and (2) a candidate retrieval module that effectively models domain-specific semantic similarity to filter noisy entities while prioritizing high-quality candidates.
@@
Variant 3 $\checkmark$ $\checkmark$ $\checkmark$ 16.83 17.13 18.71 21.87 20.54 25.09 18.69 21.36
@@
To assess the contribution of each module in AutoPCR and answer (RQ-2), we conduct ablation experiments shown in Table 4. Specifically, we evaluate the following four variants of AutoPCR. Variant 1 removes syntactic entity extraction and relies solely on bioNER for entity extraction. It assesses whether syntactic entities contribute beyond direct BioNER segmentation. Variant 2 disables HPO knowledge and retrains SapBERT (initialized from PubMedBERT) on UMLS excluding all HPO-related concepts. It simulates a scenario where the target ontology is newly introduced and unseen during model training. Variant 3 examines the necessity of domain-specific alignment in SapBERT by replacing it with PubMedBERT for embedding, without any fine-tuning on UMLS. Variant 4 evaluates the importance of entity linking by entirely removing the linking module and rely solely on candidate retrieval for concept prediction. It makes prediction only if the top candidate concept exceeds the similarity threshold $\tau_{1}$.
The full AutoPCR model overall outperforms all ablated variants across datasets and metrics, confirming the importance of each component. Variant 1 shows a clear performance decline on BIOC-GS, confirming that BioNER alone is insufficient for handling free-form text and that syntactic entity extraction is essential for improving recall in such cases. Variant 3 exhibits the most severe performance drop, demonstrating that domain-specific alignment in embedding-based retrieval is critical for accurate concept grounding. Variant 4 performs reasonably well on standardized abstracts such as GSC-2024 and ID-68, suggesting that surface-form matching alone may suffice for low-ambiguity datasets. However, it underperforms markedly on BIOC-GS, which contains noisier and more ambiguous clinical narratives, underscoring the value of LLM-based reasoning for fine-grained semantic disambiguation. Notably, Variant 2 achieves the strongest results among all variants and remains comparable to the full AutoPCR model even without HPO-specific knowledge, indicating that AutoPCR can perform inductive inference on entirely new or rapidly evolving ontologies, such as HPO.
@@
Table 5: Performance of selected methods on NCBI. AutoPCR obtains the highest accuracy with the lowest deployment time.
@@
To evaluate the generalizability and deployment efficiency of AutoPCR, we conduct an additional experiment on the NCBI dataset using the MEDIC ontology, as shown in Table 5. We define deployment time as the total time needed to prepare a method for inference on a new ontology, including ontology-specific index construction or model retraining. We compare AutoPCR with three strong and representative baselines: FastHPOCR, NCR, and PhenoTagger. PhenoTagger++ and PhenoBERT are excluded due to the lack of publicly available pipelines for adapting to a new ontology.
@@
Table 6: Performance of AutoPCR with various open-source and proprietary LLM backends.
LLM Backend BIOC-GS GSC-2024 ID-68 NCBI Avg. mF1 Avg. dF1
mF1 dF1 mF1 dF1 mF1 dF1 mF1 dF1
Open- Source Llama-3.1-8B-Instruct 70.86±0.32 71.06±0.32 79.93±0.08 79.83±0.10 78.50±0.13 78.74±0.12 69.75±0.04 65.62±0.04 74.76±0.14 73.81±0.15
Llama-3.3-70B-Instruct 72.63±0.19 72.85±0.20 82.33±0.04 82.72±0.05 79.20±0.11 79.56±0.13 71.00±0.02 68.07±0.14 76.29±0.09 75.80±0.13
Llama-4-Scout-17B-16E-Instruct 72.41±0.25 72.70±0.25 82.14±0.22 82.25±0.31 80.90±0.12 81.16±0.11 72.93±0.21 70.74±0.22 77.09±0.20 76.71±0.23
Qwen3-30B-A3B-Instruct-2507 72.60±0.03 72.81±0.03 81.71±0.01 82.03±0.00 78.73±0.00 79.07±0.00 71.36±0.02 68.84±0.00 76.10±0.01 75.69±0.01
Qwen3-Next-80B-A3B-Instruct 74.81±0.13 75.04±0.13 82.97±0.09 83.48±0.11 79.98±0.11 80.34±0.11 73.15±0.18 71.93±0.34 77.73±0.13 77.70±0.17
DeepSeek-V3.1 73.66±0.21 73.88±0.21 83.15±0.03 83.80±0.09 80.14±0.07 80.42±0.07 72.47±0.02 71.21±0.08 77.35±0.08 77.33±0.11
Kimi-K2-Instruct-0905 72.77±0.12 72.98±0.12 83.00±0.25 83.70±0.31 79.33±0.14 79.74±0.17 71.46±0.37 68.70±1.23 76.64±0.22 76.28±0.46
Proprietary GPT-4o-mini 73.95±0.11 74.18±0.11 82.71±0.09 83.04±0.10 79.36±0.10 79.63±0.11 71.54±0.05 69.05±0.07 76.89±0.09 76.47±0.10
GPT-4.1-nano 72.55±0.24 72.76±0.24 81.92±0.06 81.93±0.10 79.27±0.12 79.55±0.13 71.06±0.10 68.54±0.15 76.20±0.13 75.69±0.16
GPT-4.1-mini 73.66±0.07 73.88±0.07 82.67±0.03 82.90±0.02 79.18±0.09 79.55±0.12 71.79±0.09 68.80±0.09 76.83±0.07 76.28±0.07
GPT-4.1 74.41±0.19 74.63±0.19 83.12±0.07 83.64±0.11 79.46±0.05 79.84±0.05 72.28±0.06 69.91±0.17 77.32±0.09 77.00±0.13
@@
Table 6 summarizes the performance of AutoPCR across eleven representative non-reasoning LLM backends (Grattafiori et al., 2024; Meta, 2025; Liu et al., 2024; Team et al., 2025) under three random seeds. We have the following findings: (1) The choice of LLM family has only a modest effect on performance. Flagship models from different families, such as Llama-4-Scout, Qwen3-Next-80B, DeepSeek-V3.1, and GPT-4.1, exhibit a performance delta of less than 1.0 F1, demonstrating the robustness of AutoPCR across various LLM backends. (2) Random seeds have negligible impact. The variance in F1 remains below 0.2 for most backends except Kimi-K2 and Llama-4-Scout, confirming the stability of AutoPCR. (3) Performance generally scales with model size. Within the Llama family, F1 scores rise steadily from the 3B to the 70B and 109B (Scout) models. Similar upward trends are observed across the GPT family (from nano to mini to GPT-4.1) and the Qwen series (from 30B to 80B). The large-scale DeepSeek-V3.1 (685B) also performs strongly, ranking near the top. These results highlight the scalability of AutoPCR and its ability to benefit from increasingly capable LLMs. (4) For practical deployment, the open-source Qwen3-Next-80B offers the best overall performance with reasonable inference efficiency, while for proprietary settings, smaller GPT-mini models provide an excellent trade-off between accuracy and cost.
@@ 5 Conclusion
In this work, we have developed AutoPCR, an automated phenotype CR method by prompting. Our experiments show that AutoPCR not only achieves superior accuracy and robustness on HPO, but also exhibits strong inductive capability and generalizability to an unseen ontology, hence the name “automated”.
@@ 8 Data availability
Datasets and ontologies are sourced from Weissenbacher et al. (2023, BIOC-GS), Groza et al. (2024b, GSC-2024), Feng et al. (2022, ID-68), Luo et al. (2021, NCBI), Talapova et al. (2023, HPO), and Davis et al. (2025, MEDIC). Access links are available in our code repository.

### 02 Apr 2026 v3

@@ 4 Experiments
On BIOC-GS consisting of noisy clinical notes, we observe a clear trend across method categories: prompt-based methods outperform neural methods, which in turn surpass dictionary-based ones, consistent with prior findings (groza2024evaluation; qi2024improved; shlyk2024real). This pattern reflects their respective modeling capabilities. Prompt-based methods use LLMs with strong language understanding abilities for entity linking, enabling them to better handle free-form and lexically diverse clinical text. Neural methods, typically fine-tuned on structured, ontology-aligned corpora, struggle to generalize to such noisy input. Dictionary-based methods perform worst, as they lack semantic understanding and rely solely on exact or approximate string matching.
@@
On GSC-2024 and ID-68, which feature more standardized text with less surface-level variation in entities, AutoPCR remains highly competitive, even though the advantage of language understanding becomes less pronounced. The top-performers on these two datasets, FastHPOCR and PhenoBERT, do not bring the advantage to the other; by contrast, AutoPCR consistently ranks behind them in the second place. This may be possibly because both methods partially annotated the datasets themselves, introducing inductive bias in their favor. This is also reflected in the experimental results, where PhenoBERT achieves an outlier precision on ID-68, hinting at a bias in the dataset’s construction (more in the Discussion section). Nevertheless, AutoPCR still secures the second best, outperforming the next best baseline by approximately 3% in F1. Similarly, AutoPCR maintains a clear advantage over all other baselines except FastHPOCR on GSC-2024. This consistent lead demonstrates that AutoPCR remains robust even under deviating annotation criteria. Unlike on BIOC-GS, AutoPCR here significantly outperforms REAL, which can be credited to two novel designs: (1) a comprehensive entity extraction approach that improves recall on long, standardized text compared to LLM-based extraction and (2) a candidate retrieval module that effectively models domain-specific semantic similarity to filter noisy entities while prioritizing high-quality candidates.
@@
Variant 3a $\checkmark$ $\checkmark$ 60.49 60.85 57.05 59.09 68.58 72.67 62.04 64.20
Variant 3b $\checkmark$ $\checkmark$ 73.02 73.31 77.77 79.89 77.72 77.77 76.17 76.99
@@
Ablation study
To assess the contribution of each module in AutoPCR and answer (RQ-2), we conduct ablation experiments shown in Table 3. Specifically, we evaluate the following four variants of AutoPCR. Variant 1 removes syntactic entity extraction and relies solely on bioNER for entity extraction. It assesses whether syntactic entities contribute beyond direct BioNER segmentation. Variant 2 disables HPO knowledge and retrains SapBERT (initialized from PubMedBERT) on UMLS excluding all HPO-related concepts. It simulates a scenario where the target ontology is newly introduced and unseen during model training. Variant 3a examines the necessity of domain-specific alignment in SapBERT by replacing it with PubMedBERT for embedding, without any fine-tuning on UMLS. Variant 3b replaces SapBERT with text-embedding-ada-002, the general-purpose embedding model used by REAL, to isolate AutoPCR’s architectural contribution from its retrieval model choice. Implementation details of both variants are in Supplementary Section 1. Variant 4 evaluates the importance of entity linking by removing the linking module and relying solely on candidate retrieval for concept prediction. It makes prediction only if the top candidate concept exceeds the similarity threshold $\tau_{1}$.
The full AutoPCR model overall outperforms all ablated variants across datasets and metrics, confirming the importance of each component. Variant 1 shows a clear performance decline on BIOC-GS, confirming that BioNER alone is insufficient for handling free-form text and that syntactic entity extraction is essential for improving recall in such cases. Variant 3a exhibits the most severe performance drop, demonstrating that domain-specific alignment in embedding-based retrieval is critical for accurate concept grounding. Variant 3b, trailing AutoPCR by approximately 2 F1 points on average, corroborates this finding, while still exceeding REAL by nearly 10 F1 points, highlighting the strength of AutoPCR as a framework over REAL independent of the retrieval model. Variant 4 performs reasonably well on standardized abstracts such as GSC-2024 and ID-68, suggesting that surface-form matching alone may suffice for low-ambiguity datasets. However, it underperforms markedly on BIOC-GS, which contains noisier and more ambiguous clinical narratives, underscoring the value of LLM-based reasoning for fine-grained semantic disambiguation. Notably, Variant 2 achieves the strongest results among all variants and remains comparable to the full AutoPCR model even without HPO-specific knowledge, indicating that AutoPCR can perform inductive inference on entirely new or rapidly evolving biomedical ontologies, such as HPO.
@@
Table 4: Performance of selected methods on NCBI. AutoPCR obtains the highest accuracy with one of the lowest deployment times.
@@
To evaluate the generalizability and deployment efficiency of AutoPCR, we conduct an additional experiment on the NCBI dataset using the MEDIC ontology, as shown in Table 4. We define deployment time as the total time needed to prepare a method for inference on a new ontology, including ontology-specific index construction or model retraining. We compare AutoPCR with representative baselines: FastHPOCR, NCR, PhenoTagger, Vanilla, and REAL. PhenoTagger++ and PhenoBERT are excluded due to the lack of publicly available pipelines for adapting to a new ontology.
@@
Supplementary Table 2 summarizes the performance of AutoPCR across eleven representative non-reasoning LLM backends (grattafiori2024llama; meta2025llama; liu2024deepseek; team2025kimi) under three random seeds. We have the following findings: (1) The choice of LLM family has only a modest effect on performance. Flagship models from different families, such as Llama-4-Scout, Qwen3-Next-80B, DeepSeek-V3.1, and GPT-4.1, exhibit a performance delta of less than 1.0 F1, demonstrating the robustness of AutoPCR across various LLM backends. (2) Random seeds have negligible impact. The variance in F1 remains below 0.2 for most backends except Kimi-K2 and Llama-4-Scout, confirming the stability of AutoPCR. (3) Performance generally scales with model size. Within the Llama family, F1 scores rise steadily from the 3B to the 70B and 109B (Scout) models. Similar upward trends are observed across the GPT family (from nano to mini to GPT-4.1) and the Qwen series (from 30B to 80B). The large-scale DeepSeek-V3.1 (685B) also performs strongly, ranking near the top. These results highlight the scalability of AutoPCR and its ability to benefit from increasingly capable LLMs. (4) For practical deployment, the open-source Qwen3-Next-80B offers the best overall performance with reasonable inference efficiency, while for proprietary settings, smaller GPT-mini models provide an excellent trade-off between accuracy and cost.
5 Discussion
Through error analysis of AutoPCR on ID-68, we identified several annotation issues that likely introduce spurious errors. First, ID-68, as annotated by PhenoBERT’s team, contains mentions linked to hypernyms rather than exact ontology matches. This occurs because PhenoBERT accepts ancestral relationships as correct predictions when an exact match is absent. There are fifteen such cases where PhenoBERT correctly predicts an ancestral relationship that AutoPCR does not (e.g., “raised choline levels in the white matter” $\rightarrow$ HP:0002500 Abnormal cerebral white matter morphology). While this criterion may be useful in certain contexts, it diverges from the standard task definition shared by our study and other competitive baselines, which leads to an ambiguous annotation standard. If every term covered by the ontology’s hierarchy is a potential candidate, the task essentially shifts into a classical hierarchical classification problem. Conversely, if only a subset is chosen, the labels become heavily influenced by annotator-idiosyncratic choices, making the problem more suitable for supervised learning rather than the zero-shot setting we address. Second, ID-68 was annotated using an outdated version of HPO (2019-09-06), causing several concept labels to differ from those in the 2024-02-08 version adopted by us and most recent methods FastHPOCR and PhenoTagger++. For example, five mentions of “thin corpus callosum” are linked to HP:0002079 Hypoplasia of the corpus callosum, whereas HP:0033725 Thin corpus callosum exists as a distinct concept in the newer HPO. Similar issues affect seventeen instances of “hypotonia” and three of “high arched palate”. Third, several annotations are inconsistent with ID-68’s stated convention of preferring longest-span matches. For example, “poor attention and hyperactivity” is annotated as separate entities rather than being linked to HP:0007018 Attention deficit hyperactivity disorder, and “Arnold Chiari malformation type I” is labeled as “Arnold Chiari malformation” instead of HP:0007099 Chiari type I malformation. These findings highlight the potential for a refined version of ID-68, similar to the transition from GSC+ to GSC-2024, to better support future benchmarking.
Apart from defective ID-68 benchmarking, AutoPCR has several other limitations as an intelligent system aimed at large-scale deployment. First, the current benchmarks are relatively small and domain-specific. Performance on larger corpora such as radiology and pathology reports remains to be validated. Additionally, AutoPCR does not explicitly handle clinical uncertainty, treating mentions such as “possible” or “rule out” identically to definitive ones, which may introduce false positives in settings where distinguishing confirmed from suspected phenotypes is critical. Furthermore, its three-stage pipeline introduces inherent sequential dependencies between stages, limiting throughput at scale. System-level optimizations such as pipeline parallelism may be explored to improve efficiency. These directions are left for future work.
@@ 6 Conclusion
In this work, we have developed AutoPCR, a prompting-based automated phenotype CR method that decomposes the task into entity extraction, candidate concept retrieval, and LLM-based entity linking. Experiments show that AutoPCR achieves superior average accuracy and robustness across benchmarks, strong inductive capability to handle HPO updates, and efficient generalizability to a new ontology within minutes, hence the name “automated”.
@@ 9 Data availability
Datasets and ontologies are sourced from weissenbacher2023phenoid, groza2024fasthpocr, feng2022phenobert, luo2021phenotagger, gargano2024mode, and davis2025comparative. Access links are available in our code repository.